OpenStudy (magbak):

PLEASE HELP ME!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! I WILL AWARD MEDAL!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! c=-5/2 c=7/2 c=4 c=7

OpenStudy (magbak):

\[25^(2c)=\sqrt5^(4c-10)\]

OpenStudy (magbak):

@primeralph @Jamierox4ev3r @Callisto @Hero

OpenStudy (magbak):

Which of the following is the solution to the equation

OpenStudy (primeralph):

Can you try to input the question better?

OpenStudy (magbak):

Ok

OpenStudy (magbak):

\[25^(2c)=\sqrt5^(4c-10)\]

OpenStudy (magbak):

25^(2c)=sqrt(5)^(4c-10)

OpenStudy (primeralph):

Same thing.

OpenStudy (magbak):

This is how it is given to me

OpenStudy (magbak):

That is not a option.

OpenStudy (magbak):

Even if you simplify that you get -5

OpenStudy (magbak):

@primeralph

OpenStudy (primeralph):

Solve for c.

OpenStudy (magbak):

It is -5

OpenStudy (magbak):

See that is not an answer

OpenStudy (primeralph):

Tag someone else to see.

OpenStudy (magbak):

Ok

OpenStudy (magbak):

@Hero @Callisto @ash2326

OpenStudy (magbak):

@Luigi0210 @tkhunny

OpenStudy (magbak):

@bahrom7893

OpenStudy (magbak):

@nincompoop

OpenStudy (bahrom7893):

\[25^{2c}=\sqrt{5}^{4c-10}\]?

OpenStudy (magbak):

yes

OpenStudy (luigi0210):

Change the bases then, do you know how?

OpenStudy (bahrom7893):

okay, i'll give you one hint: \[25=5^2\] \[\sqrt{5}=5^{1/2}\]

OpenStudy (magbak):

What is that supposed to mean.

OpenStudy (luigi0210):

It means change them so they have the same base

OpenStudy (bahrom7893):

\[25^{2c}=\sqrt{5}^{4c-10}\]it means rewrite this so that there are only 5s on the bottom.

OpenStudy (magbak):

Oh

OpenStudy (magbak):

It is A. Correct me if I am wrong

OpenStudy (bahrom7893):

no guessing, show us the solution

OpenStudy (magbak):

Well I squared both sides so that is what I got. @campbell_st

OpenStudy (magbak):

@bahrom7893

OpenStudy (magbak):

PLease @bahrom7893 @campbell_st @Jamierox4ev3r @ash2326 @Callisto @primeralph @Preetha

OpenStudy (magbak):

Help me

OpenStudy (luigi0210):

Should I keep going?

OpenStudy (magbak):

PLease

OpenStudy (luigi0210):

Continuing off of what bah said, change the bases

OpenStudy (magbak):

HOW

OpenStudy (luigi0210):

Okay.. you know that: \[25=5^2\] and \[\sqrt{5}=5^{1/2}\] right?

OpenStudy (magbak):

Yes

OpenStudy (magbak):

Now what

OpenStudy (luigi0210):

Change the bases in your equation: \[25^{2c}=\sqrt{25}^{4c-10}\] is the same as: \[5^{2(2c)}=5^{1/2(4c-10)}\] When you have same bases you can just set the exponents equal to each other: 2(2c)=1/2(4c-10)

OpenStudy (magbak):

I AM LOST

OpenStudy (magbak):

@agent0smith

OpenStudy (primeralph):

8c = 4c-10.

OpenStudy (luigi0210):

I really don't know how to make it any more simpler..

OpenStudy (magbak):

OK @primeralph

OpenStudy (magbak):

It is A like I SAID THE FIRST TIME

OpenStudy (primeralph):

Yeah.

OpenStudy (magbak):

THNX @primeralph

OpenStudy (magbak):

MEPS WHAT YOU LOOKING AT KEEP MOVING

OpenStudy (anonymous):

Make me

OpenStudy (primeralph):

Meps.

OpenStudy (anonymous):

yo prime... its the latex

OpenStudy (magbak):

Just joking relax

OpenStudy (magbak):

I like your picture

OpenStudy (anonymous):

come at me bro

OpenStudy (campbell_st):

ok... all you need to do is have both sides mof the equation with the same base... which will be 5 and apply the power of a power law \[(x^a)^b = x^{a \times b}\] so you'll have \[(5^2)^{2c} = (5^{\frac{1}{2}})^{4c - 10}\] apply the power of a power law \[5^{2 \times 2c} = 5 ^{\frac{1}{2} \times (4c -10)}\] so you are looking at \[5^{4c}= 5^{2c - 5}\] since the based are the same you can equate the powers so you need to solve 4c = 2c - 5 for c hope this makes sense